Q1. Find the density function of Y = ex if X is normally distributed with mean μ and standard

deviation σ.


Q2. Find the density function of Y = βX(1/α) where α > 0, β > 0 if X is exponentially distributed

with mean μ = 1.



1
Expert's answer
2021-12-09T14:41:08-0500

1.

density for Normal distribution:


"f(x)=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{1}{2}(\frac{x-\mu}{\sigma})^2}"


 cumulative distribution function:


"F_Y(y)=\int^{lny}_{-\infin}f(x)dx"


To find the density, differentiate. Using the Fundamental Theorem of Calculus:

"f_Y(y)=\frac{1}{y}f(lny)"


"f_Y(y)=\frac{1}{y}\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{1}{2}(\frac{lny-\mu}{\sigma})^2}"


2.

"f(x)=\lambda e^{-\lambda x}=e^{-x}"

"X=(Y/\beta)^{\alpha}"


"f_Y(y)=f_X((Y/\beta)^{\alpha})\frac{d}{dy}((Y/\beta)^{\alpha})=\alpha e^{-(y/\beta)^{\alpha}}y^{\alpha -1}/\beta^{\alpha}"